Kerron Company is presented with the following two mutually exclusive projects. The required return for both projects is 15 percent. Year Project M Project N 0 -$125,000 -$310,000 1 57,000 135,000 2 64,000 161,000 3 59,000 129,000 4 34,000 92,000 Required: (a) What is the IRR for each project? (Do not round intermediate calculations. Enter your answer as a percentage rounded to 2 decimal places (e.g., 32.16).) IRR Project M% Project N % (b) What is the NPV for each project? (Do not round intermediate calculations. Round your answers to 2 decimal places (e.g., 32.16).) NPV Project M $ _________ Project N $ _________
Answer #1

Hi,

Please find the detailed answer as follows:

Part A:

Project M :

To calculate IRR, you need to put the value of NPV as 0 and solve for r as follows:

NPV = 0 = -125000 + 57000/(1+r)^1 + 64000/(1+r)^2 + 59000/(1+r)^3 + 34000/(1+r)^4

Solving for r we get IRR as 27.70%

IRR = 27.70%

------

Project N:

NPV = 0 = -310000 + 135000/(1+r)^1 + 161000/(1+r)^2 + 129000/(1+r)^3 + 92000/(1+r)^4

Solving for r we get IRR as 25.91%

IRR = 25.91%

IRR Table:

IRR
Project M 27.70%
Project N 25.91%

-------

Part B: NPV

Project M = -125000 + 57000/(1+15%)^1 + 64000/(1+15%)^2 + 59000/(1+15%)^3 + 34000/(1+15%)^4 = $31191.48

----

Project N = -310000 + 135000/(1+15%)^1 + 161000/(1+15%)^2 + 129000/(1+15%)^3 + 92000/(1+15%)^4 = $66551.33

-----

NPV Table:

NPV
Project M 31191.48
Project N 66551.33

Thanks.
Answer #2

Calculation of IRR Calculation of NPV Project M Project N PVF(15%) Amount 0 Year Project VM 0 $(125,000.00) $(310,000.00) 1 $